# Attribute-driven Disentangled Representation Learning for Multimodal Recommendation

### v1

@@ 4.4. Interpretability Study (RQ3)
4.5. Controllability Study (RQ3)
One of the primary targets of our method is to improve the controllability of recommendation systems. In this section, we assess the controllability of AD-DRL by removing the representation chunks of factors associated with price and popularity. As shown in Figure 4(a), for the attribute of price, we selected 100 users who always purchase inexpensive products based on their purchase records. ”w/ price” indicates the recommendation results of the AD-DRL model utilizing all factors, where the model tends to recommend inexpensive products to users. However, when we removed the chunk associated with the price (denoted as ”w/o price” in Figure 4(a)), we found that the recommended products from AD-DRL were more diverse in terms of price. This underscores AD-DRL’s capability for attribute-level controllability. It is with noting that after removing the price attribute, the products recommended by AD-DRL still tend to be inexpensive. This could be attributed to the retained chunk associated with the brand, and there are correlations between brand and price. For example, brand reputation plays a significant role in pricing. Some high-end brands offer products at a premium price, while others provide more affordable options. Similar results can be seen in another group of users who consistently purchase expensive products. As a comparison, due to the lack of a clear definition for each factor in previous work, the results of removing the factors associated with specific attributes are unpredictable.
Figure 4. Statistics of the recommendation results while AD-DRL removes the chunk of price factor or popularity factor. ##\##counts refer to the number of items with different attribute values.
4.6. Ablation Study (RQ4)
@@
[…]
@@ 5. Conclusion
Existing disentangled representation learning methods in recommender systems disentangle the underlying factors behind user-item interaction data in an unsupervised manner, leading to limit interpretability and controllability of recommendation. This paper proposes an attribute-driven disentangled representation learning method, termed AD-DRL, which disentangles attribute factors in various multimodal features. For obtaining robust and independent representations, the proposed AD-DRL enables disentangling within each modality feature and ensures consistency of representations across modalities at the attribute level. Moreover, it leverages the intrinsic relationships between items sharing the same attribute value. Experimental results demonstrate the superiority of our method AD-DRL compared to existing methods and showcase its capability in terms of interpretability and controllability.

### v2

@@ 4.4. Interpretability Study
4.5. Controllability Study
(a) Users who prefer low-price level items.
(b) Users who prefer high-price level items.
(c) Users who prefer low-popular level items.
(d) Users who prefer high-popular level items.
Figure 5. The proportion of items with different attribute values within the AD-DRL’s recommendations when $\xi$ takes different values in Equation 14.
In this section, we evaluate the controllability of AD-DRL by manipulating user preferences for certain attributes. Specifically, we change the user $u$’s preference score for a specific attribute $a$ in Equation 11 to assess if such change yields the anticipated variation in recommendation results. We modify the formula for $s_{u,i}$ in Equation 11 as follows:
(14) $s_{u,i}=\xi*s_{u,i,a}+\sum_{k\neq a}s_{u,i,k},$
where $\xi$ represents the scaling factor for $s_{u,i,a}$, used to adjust the impact of attribute $a$ on $s_{u,i}$. For a specific attribute (in this section, we selected price and popularity for study), we systematically vary $\xi$ through values of 2, 1, 0.5, 0, and -1. This process allows us to meticulously examine the resultant shifts in our method’s recommendation outputs, with findings illustrated in Figure 5.
We conduct a detailed analysis using the price attribute as an example. As shown in Figure 5a, we select 100 users from Baby dataset who always purchase inexpensive items based on their purchase records. Various colors in Figure 5a denote distinct price levels, showcasing the distribution of items recommended by AD-DRL across these levels for a given value of $\xi$. By comparing the recommendations made by AD-DRL under different values of $\xi$, we have the following observations:
  • When $\xi=1$, that is, using the original AD-DRL method, it can be seen that AD-DRL is capable of capturing the preference of these users for inexpensive items and recommends more affordable items to them.
  • When $\xi=2$, meaning we have increased the impact of price on the user preference score, it can be observed that AD-DRL tends to recommend more inexpensive items to these users.
  • When $\xi=0.5$ or 00, meaning we reduce or even eliminate the impact of price on the user preference score, we find that the recommended items from AD-DRL are more diverse in terms of price. It is worth noting that even when $\xi=0$, the items recommended by AD-DRL still tend to be inexpensive. This could be attributed to the retained chunk associated with the brand, and there are correlations between brand and price. For example, some high-end brands offer products at a premium price, while others provide more affordable options.
  • More interestingly, when $\xi=-1$, meaning we have AD-DRL recommend items that are opposite to the users’ price preferences, it can be observed that AD-DRL indeed recommends some more expensive items.
The above results reveal that adjusting the value of $\xi$ enables us to align AD-DRL’s recommendation results with our expectations, thereby illustrating the controllability of AD-DRL. Similar results can be seen in another group of users who prefer high-price level items.
4.6. Ablation Study
@@
Table 3 shows the experimental results for all variants. From the results, we have the following observations:
  • Using either high-level or low-level disentangled representation learning independently can significantly improve the performance. This indicates that both disentangled representation learning within modalities and between modalities have a positive effect.
  • The AD-DRL combines high-level and low-level disentangled representation learning modules, resulting in significantly improved performance compared to the devised two variants. This demonstrates the necessity and rationality of combining the attribute and attribute value levels.
  • Delving deeper into the high-level disentangled representation learning module, removing the intra-modality disentanglement module results in a greater decrease in model performance compared to removing the inter-modality disentanglement module. This indicates that intra-modality disentanglement is the fundamental aspect of disentangled representation learning, while inter-modality disentanglement can further enhance model performance.
@@ 5. Conclusion
In this paper, we highlight the limitations of existing disentangled representation learning techniques in recommender systems. The current methods disentangle the underlying factors behind user-item interactions in an unsupervised manner, leading to limited interpretability and controllability. To overcome this limitation, we propose an attribute-driven disentangled representation learning method, termed AD-DRL, which disentangles attribute factors in various multimodal features. More specifically, the proposed AD-DRL enables disentangling within each modality feature and ensures consistency of representations across modalities at the attribute level. Moreover, it leverages the intrinsic relationships between items sharing the same attribute value. Experimental results demonstrate the superiority of our proposed AD-DRL and showcase its capability in terms of interpretability and controllability.